OpenStudy (anonymous):

Two players kick a soccer ball at the same time. If one player applies a force of 100N (N 25 degrees W ) and the 250g ball experiences an acceleration of 200m/s^2 (W 15 degrees S), determine the magnitude and direction of the force applied by the second player.

OpenStudy (anonymous):

I just wanna know the algorithm of this problem. If i can't solve it, i'll ask for the mathematical steps to get the answer.

OpenStudy (anonymous):



OpenStudy (anonymous):

Find the components of F1 and a x component of F1/m + x component of F2/m = x component of a y component of F1/m + y component of F2/m = y component of a

OpenStudy (anonymous):

remember to convert 250 g to kg before you start

OpenStudy (anonymous):

can't you just convert a to F2 using the mass in kg So F2 = ma F2 = 50N Would that get the same answer?

OpenStudy (anonymous):

direction?

OpenStudy (anonymous):

direction?

OpenStudy (anonymous):

"determine the magnitude and direction of the force"

OpenStudy (anonymous):

Also, your way, whatever it is, totally neglects the 100N force..

OpenStudy (anonymous):

Wait, so you have to divide each vector by the mass?

OpenStudy (anonymous):

And no that was just the gist of my way. I actually converted the acceleration to force using the mass Then find the resultant now that both the vectors are the same unit

OpenStudy (anonymous):

if you don't want to divide by mass, you don't have to... totally fine to solve it as x component of F1 + x component of F2 = x component of m*a y component of F1 + y component of F2 = y component of m*a

OpenStudy (anonymous):

your choice:)

OpenStudy (anonymous):

I am reallylost. :P

OpenStudy (anonymous):

ask a question or something, this is pretty straight forward, so I don't know where you're getting lost...

OpenStudy (anonymous):

can you do that and tell me what you get, I have the answers so I can see if your ways is correct. Just do it on paper since I know what you're talking about. And since it's much faster

OpenStudy (anonymous):

F/m is an acceleration ... do you agree? so the x component of F1/m (the acceleration caused by Force 1) PLUS the x component of F2/m (the acceleration caused by Force 2) must equal the net x acceleration.... (which we are given)

OpenStudy (anonymous):

Welps, I didn't do the question right. I understand what you're saying :D Thanks, you have a new fan:)